I have received two conflicting answers.
One tutor says 78.03, the other says 78.08

All calculations should NOT be rounded, but the final answer should be (as noted in directions).
What is the correct answer?

| Question 5 2 pts Lion Inc., has sales of $2550, total assets of $1021, and a debt-equity ratio of 0.7. If its return on equity is 13%. What is Lions Net Income? (Round final answer to 2 decimal places. Do not round intermediate calculations). Topic: Financial Ratios (DuPont Model) 78.08
Answer #1

Equity = Assets / (Debt + Equity)

            = 1,021 / (0.7 + 1)

            = 1,021 / 1.7

            = 600.58

DuPont model:

Return on equity = Net income ratio × Asset turnover ratio × Equity multiplier

0.13 = (Net income / 2,550) × (Sales / Assets) × (Assets / Equity)   

0.13 = (Net income / 2,550) × (2550 / 1021) × (1021 / 600.58)

0.13 = Net income / 600.58

Net income = 0.13 × 600.58

                    = 78.0754           

                    = 78.08 by rounding (Answer)